# Approximate Quadratization of High-Order Hamiltonians for Combinatorial Quantum Optimization

### v1

@@ III.1 Method
We quadratize HUBOs with a hypergraph clique expansion [55]. The price paid to keep the qubit count constant is an approximation of the original problem. We treat a fourth-order Hamiltonian as a hypergraph $\mathcal{G}=(\mathcal{E},V)$ by identifying each four-local $Z_{j}Z_{k}Z_{h}Z_{l}$ term with a hyperedge connecting nodes $e=(j,k,h,l)\in\mathcal{E}$. To quadratize $H_{4}$, we create a weighted graph $G=(E,V)$, resulting in a quadratic Hamiltonian $H_{2}$. We add an edge to $E$ for each pair of nodes $(j,k)$ within a hyperedge $e$, as shown in Fig. 2(c). Therefore, each hyperedge in $\mathcal{E}$ generates a four-clique in $G$. Since a given pair of nodes $(j,k)$ might belong to multiple hyperedges of $\mathcal{E}$, the weight $w_{2}^{\text{opt}}(j,k)$ we assign to it minimizes the $l_{2}$-norm, i.e. the sum of squared differences, with respect to the weights $w_{4}(e)$ of all hyperedges $e$ that contain $(j,k)$. The $l_{2}$-norm is chosen according to the quadratization given in the literature [55]. Concretely, for each resulting edge $(j,k)\in E$, we consider the set of hyperedges $\mathcal{E}_{(j,k)}$ that contain the edge $(j,k)$ and choose
$w_{2}^{\text{opt}}(j,k)=\arg\min_{w_{2}(j,k)}\sum_{e\in\mathcal{E}_{(j,k)}}\Big{(}w_{2}(j,k)-w_{4}(e)\Big{)}^{2}.$ (7)
Figure 2: Construction of shallow Ansatze. (a) and (b) show how a four-local gate $\exp(-i\gamma ZZZZ)$ and a two-local gate $\exp(-i\gamma ZZ)$, respectively, decompose into $R_{Z}$ and $CX$ gates. (c) The hypergraph clique expansion generates a 2D graph that is easier to implement on quantum hardware. The graph is generated such that each pair of nodes in the original hypergraph has a corresponding edge, and the weight of this edge is given by Eq. (7). Here, the edge $(0,1)$ has weight $w_{c}=(w_{a}+w_{b})/2$, whereas the other edges preserve the weights of the single hyperedge they correspondingly belong to. (d) Circuit implementation of the resulting graph with layers of SWAP gates. To avoid hardware noise, we skip implementing the last two SWAP layers, at the cost losing some correlation with the original problem.
We use the resulting quadratic Hamiltonian
$H_{2}=\sum_{(j,k)\in E}w_{2}^{\text{opt}}(j,k)Z_{j}Z_{k}$ (8)
@@
|$\ket{\psi_{2}(\bm{\beta},\bm{\gamma})}=\prod_{q=1}^{p}e^{-i\beta_{q}H_{M}}e^{-i\gamma_{q}H_{2}}\ket{+}^{\otimes n},$ (9)
we minimize the energy $\langle\psi_{2}(\bm{\beta},\bm{\gamma})|H_{4}|\psi_{2}(\bm{\beta},\bm{\gamma})\rangle$.
@@
Our approach introduces an inconsistency between the cost operator and the operator that creates the Ansatz. Therefore, the convergence guarantee of a typical QAOA in the limit $p\to\infty$ is lost. Instead, we are left with a heuristic approach based on an Ansatz circuit that retains only some of the structure of $H_{4}$, but which considerably reduces the circuit depth. The exact depth reduction depends on the density of the original problem. The largest reduction occurs for dense Hamiltonians, such as $H_{4}^{\text{full}}$, for which the reduction scales as $O(n^{2})$, see Fig. 1. For LABS, which has $O(n^{3})$ terms, the reduction factor is $O(n)$.
III.2 Expected performance of quadratization
We now assess the quadratization by creating a scatter plot of the energies of the quartic Hamiltonian $\langle x|H_{4}|x\rangle$ against the energies of the corresponding quadratic Hamiltonian $\langle x|H_{2}|x\rangle$, for all basis states $x$. We perform this assessment on the $H_{4}^{\text{full}}$ and LABS Hamiltonians with 12 decision variables. Based on this analysis, we assess whether it is promising to look at similar problem instances for larger systems, since trends in expected performance change smoothly with system size [56].
We do not observe any correlation between the energy calculated with $H_{4}^{\text{full}}$ and with its quadratized counterpart, see Fig. 3(a). Indeed, both energy distributions are normal and uncorrelated. This implies that all of the problem structure of $H_{4}^{\text{full}}$ is lost in the quadratization.
However, the quadratization of LABS, discussed in App. C, retains some of the problem structure, see Fig. 3(b). This results from a combined effect of the distribution of edge weights and the hyperedge structure of LABS. LABS has all weights equal to either 2 for four-local terms, or 1 for two-local terms. This results in a quadratized QUBO with almost all equal edge weights. As a result, all basis states with Hamming weight $0\leq h\leq n$ and $n-h$ have almost the same QUBO energy, defining the energies of the vertical clusters in Fig. 3(b). More details on the impact of the problem structure and the edge weight distribution on the scatter plots are given in App. D. There, we also consider Hamiltonians with different structures and hyperedge weight distributions.
Figure 3: Structure retention in the quadratization. The black dots show the energy of different bitstrings with respect to the quartic Hamiltonian versus the energies with respect to the quadratized Hamiltonian. The curves show the Cummulative Distribution Function (CDF) of the quartic energies of uniform samples (pink), and samples from running QAOA with the quadratized Ansatz (purple), and the standard Ansatz (blue). The horizontal lines show the corresponding mean energies of these distributions. Simulations were done using 10 000 shots resulting in shot noise with a variance of 0.32 for $H_{4}^{\text{full}}$ and 0.18 for LABS.
Next, we run a depth-one QAOA using both the quartic and quadratic Hamiltonians to generate the Ansatz circuit. To optimize the parameters, we first run an exhaustive search in the space $\beta,\gamma\in[0,\pi/2]$, with a resolution of $60$ for both parameters, for a system of 6 qubits. For larger systems, optima are found through successive initialization of the parameters using the optimal values found for a system with one qubit less, see Fig. 4. Details on the parameter optimization procedure are in App. E.1.
In the $H_{4}^{\text{full}}$ case, we find $(\beta,\gamma)=(1.319,0.046)$ for the standard Ansatz and $(\beta,\gamma)=(1.481,1.565)$ for the quadratized Ansatz. We confirm through a t-test with $97.7\%$ confidence that the quadratized Ansatz outputs samples that produce the same energy expectation as uniformly sampled solutions, see Fig. 3. In our tests, we ran the circuits 10 times with 10 000 shots each. This confirms that for this Hamiltonian structure and normal hyperedge weight distribution, the QUBO Ansatz does not preserve the structure of the original $H_{4}^{\text{full}}$ problem.
For LABS, the optimized parameters $(\beta,\gamma)=(1.416,0.028)$ for the standard Ansatz generate better samples than random sampling. The QUBO Ansatz, with optimal parameters $(\beta,\gamma)=(1.480,0.168)$, yields a worse outcome distribution (mean lowest energy of $-3.338$) than using the standard Ansatz (mean lowest energy of $-9.852$), see optimization landscapes in Fig. 4. However, in this case, the QUBO Ansatz performs statistically better than random sampling, indicating that it preserves some of the structure of LABS. Crucially, the quadratic Ansatz has $O(n)$ times less two-qubit gates than the quartic Ansatz. Therefore, the quadratic Ansatz is expected to outperform the standard one under noisy conditions.
Figure 4: Optimization landscapes of depth-one QAOA for LABS on 12 qubits when using a LABS Ansatz in panels (a) and (c), and using the quadratized Ansatz in panels (b) and (d). Panels (c) and (d) show in more detail the areas in the red squares in panels (a) and (b), respectively. Simulations were done using a statevector simulator and are thus exact.
@@ III.3 Noisy simulations
The previous section showed that our proposed clique projection yields, in the noiseless case, an improved energy over random sampling, but not over the original quartic Ansatz. We now study this quadratization for QAOA depths $p=1$ and $2$ for LABS problems with sizes from $n=8$ to $16$ by simulating the corresponding quantum circuits with and without noise. First, in the absence of noise, as $p$ increases, we observe a decrease in energy for both the quartic and the quadratic Ansatze, see the blue and the purple curves in Figs. 5(a) and (d), respectively. By comparing corresponding pairs of lines from these two subfigures, we see a maximal relative energy improvement brought by adding a QAOA layer of only $13.52\%$ for the LABS Ansatz and $4.78\%$ for the QUBO Ansatz relative to the global minimum value ($-8,-28,-48$ for $8,12$, and $16$ qubits, respectively). This is in part due to the limited performance of a depth-two QAOA, but also due to the more difficult parameter training, which heavily relies on a suitable initialization [57].
@@
Figure 5: Evolution of expected performance of QAOA for LABS versus problem size, using the standard Ansatz (blue) and the quadratized Ansatz (purple) compared to random sampling (pink). Panels (a-c) correspond to depth-one QAOA with depolarizing noise strengths of 00 (i.e. noiseless), $0.001$, and $0.005$, respectively. Panels (d-f) show a depth-two QAOA with the same depolarizing noise strengths. The numbers in panels (a) and (d) indicate the number of two-qubit gates in the Ansatz. Each data point is the average of 10 circuit runs and the error bars represent shot noise over 10 000 samples.
We test whether a quadratic Ansatz has a better performance than a quartic one in the presence of noise by setting up a noisy simulation in which we apply a depolarizing channel with strength $\lambda$ after each hardware-native entangling gate. At $\lambda=0.001$, depth-one QAOA with a quartic Ansatz outperforms QAOA with a quadratic Ansatz, see Fig. 5(b). However, for problems with more than 13 decision variables, the average energy of the samples drawn from the quartic Ansatz starts increasing. This indicates that the energy gain of a more accurate Hamiltonian is eroded by the noise. This also explains why the depth-two quartic Ansatz, blue line in Fig. 5(e), reverses its downward trend in energy faster than the depth-one, at $n=10$ instead of $13$. Indeed, depth-two QAOA has twice the gate count of depth-one QAOA. For the 16 qubit systems, depth-two QAOA becomes indistinguishable from random sampling. Extrapolating the trend in Fig. 5(b) we conclude that depth-one QAOA has the same performance as random sampling for problems with more than $30-40$ decision variables.
The quadratized Ansatz presents more promising scaling prospects, making it a viable option for noisy quantum computers. In contrast to the full LABS Ansatz, the quadratic Ansatz preserves the decreasing energy trend with system size even in the presence of depolarizing noise strength of $\lambda=0.001$, see purple lines in Figs. 5(b) and (e). Therefore, as the system size increases, the performance gain from the quadratic Ansatz is strong enough to balance the added noise. This feature persists, albeit diminished, even for a stronger noise strength $\lambda=0.005$, where some signal can still be detected when using the QUBO Ansatz compared to the original Ansatz for $p=1$, see Figs. 5(c) and (f). However, at $p=2$, noise completely cancels out the signal for both the full LABS Ansatz and the quadratized Ansatz for all problem sizes tested.
These data show the benefits of quadratizing the Ansatz when working with hardware with small to intermediate noise. In particular, no performance over random sampling is expected from either Ansatz at the current hardware noise levels of $\lambda>0.01$ [58, 48]. However, our work indicates at which noise level one expects the quadratized Ansatz to outperform its full counterpart.
@@ IV Hardware-native circuits for QUBOs
Figure 6: Performance of QAOA as a function of depth and the number of SWAP layers. Each data point is the average of three RR3 graphs that require nine SWAP layers. The error bars show the standard deviation. (a) Approximation ratio of QAOA computed with a MPS evaluation of the energy. (b) Measured approximation ratios from 50 000 samples from ibm_fez. The horizontal dashed line is the $58.3\%$ average approximation ratio reported in Ref. [59]. (c) Value of the CVaR quantile $\alpha$ required to make the approximation ratio of the ibm_fez samples in (b) match those of the MPS energy evaluation in (a). The dashed lines show a simple theoretical expectation $\alpha_{\text{th}}$ based on the two-qubit gate fidelity.
We now study an approximate hardware-friendly compilation method of Ansatz circuits for QUBOs. The standard approach to route an arbitrary QAOA circuit utilizes a predetermined network of SWAP gates tailored to the qubit connectivity [60, 53, 41, 61]. We use a line of qubits to benefit from two-qubit gate cancellations, even though the hardware connectivity has a heavy-hex structure [60]. To fully implement $e^{-i\gamma H_{C}}$, we apply $k_{\text{max}}$ alternating layers of even and odd SWAP layers [^2] A SWAP layer is a depth-one circuit which applies SWAP gates to pairs of neighboring qubits, and it is called even if the first pair is $(0,1)$ and odd if it is $(1,2)$., see Fig. 2(d). Full connectivity is reached after $n-2$ SWAP layers. The SWAP network thus allows us to derive an approximate Ansatz by including only $k<k_{\text{max}}$ layers when implementing $e^{-i\gamma H_{C}}$. Hence, we implement only the time evolution of a subset $H_{C}(k)$ of the terms in $H_{C}$. The truncated $H_{C}(k)$ is based on the subgraph $G_{k}=(E_{k},V)\subseteq G=(E,V)$. Details on the transpilation are given in App. B.2. Importantly, the size $|E_{k}|$ depends on the initial mapping of the decision variables to the qubits. Ref. [53] proposes a method to optimize the initial mapping by using a SAT solver. Once the Ansatz is created, we optimize the QAOA parameters by minimizing the energy of the original cost operator
$\min_{{\bm{\beta}}_{k},{\bm{\gamma}_{k}}}\langle\psi_{k}({\bm{\beta}}_{k},{\bm{\gamma}}_{k})|H_{C}|\psi_{k}({\bm{\beta}}_{k},{\bm{\gamma}_{k}})\rangle$ (10)
@@
|$\ket{\psi_{k}({\bm{\beta}}_{k},{\bm{\gamma}_{k}})}=\prod_{j=1}^{p}e^{-i\beta_{j,k}H_{M}}e^{-i\gamma_{j,k}H_{C}(k)}\ket{+}^{\otimes n}.$ (11)
Our approach is similar to the time-block Ansatz design presented in Ref. [41]. The main distinction is that a time-block Ansatz assigns different, independently optimized parameters to the edges of $H_{C}(k)$. Concretely, they replace the values $\gamma_{j,k}$ with corresponding vectors $\{\gamma_{j,k}^{e},\forall e\in E\}$.
We benchmark the truncated Ansatz using Max-Cut problems [15] for all $k\in[0,k_{\max}]$. We thus aim to split the set $V$ of $n$ nodes of a graph $G=(E,V)$ into two such that the sum of the edges connecting the two subsets is maximized. Formally, a variable $x_{i}$ is assigned to each node $i\in V$ to indicate which side of the cut the node is on. Solving Max-Cut is then equivalent to finding the binary string $x$ that maximizes
$f(x)=\sum_{(j,k)\in E}\Big{(}x_{i}(1-x_{j})+x_{j}(1-x_{i})\Big{)}.$ (12)
Mapping $f(x)$ to a Hamiltonian [40] results in the cost operator
$H_{\text{Max-Cut}}=\frac{1}{4}\sum_{(j,k)\in E}Z_{i}Z_{j},$ (13)

### v2

@@ III.1 Methods
In the hypergraph clique expansion [suppakitpaisarn2024utilizing] we treat a high-order Hamiltonian $H_{C}=\sum_{i=0}^{K-1}\omega_{i}P_{i}$ as a hypergraph $\mathcal{G}=(\mathcal{E},V)$. Here, $\omega_{i}$ is the coefficient of the Pauli $P_{i}$. We associate each Pauli term $P_{i}=Z_{i_{1}}Z_{i_{2}}...Z_{i_{k}}$ with a hyperedge $e=(i_{1},i_{2},...,i_{k})\in\mathcal{E}$. To quadratize $H_{C}$, we create a weighted graph $G=(E,V)$, resulting in a quadratic Hamiltonian $H_{2}=\sum_{ij}\omega_{ij}Z_{i}Z_{j}$. We add an edge to $E$ for each pair of nodes $(i,j)$ within a hyperedge $e$, as shown in Fig. 2(c). Therefore, each hyperedge $e$ with $k$ nodes in $\mathcal{E}$ generates a $k$-clique in $G$ with edge-set $E(e)$. Since a given pair of nodes $(i,j)$ might belong to multiple hyperedges of $\mathcal{E}$, the weight $w_{ij}^{\text{opt}}$ we assign to it minimizes the $l_{2}$-norm, i.e. the sum of squared differences, with respect to the weights $w(e)$ of all hyperedges $e$ that contain $(i,j)$. The $l_{2}$-norm is chosen according to the quadratization given in the literature [suppakitpaisarn2024utilizing]. In particular, we choose $\omega_{ij}$ to minimize
$f_{2}(\bm{\omega})=\sum_{e\in\mathcal{E}}\sum_{(i,j)\in E(e)}\alpha_{e}[\omega_{ij}-\omega_{e}]^{2}.$ (7)
Here, $\omega_{e}$ is the weight of hyperedge $e$. The weight $\alpha_{e}=\omega_{e}/\sum_{e\in\mathcal{E}}\omega_{e}$ prevents low-weight edges from having a disproportionate impact on the quadratization. We use the resulting quadratic Hamiltonian
$H_{2}=\sum_{(i,j)\in E}\omega_{ij}^{\text{opt}}Z_{i}Z_{j}$ (8)
@@
[…]
we minimize the energy $\langle\psi_{2}(\bm{\beta},\bm{\gamma})|H_{C}|\psi_{2}(\bm{\beta},\bm{\gamma})\rangle$.
Figure 2: Construction of shallow Ansatze. (a) and (b) show how a four-local gate $\exp(-i\gamma ZZZZ)$ and a two-local gate $\exp(-i\gamma ZZ)$, respectively, decompose into $R_{Z}$ and CXCX gates. (c) The hypergraph clique expansion generates a 2D graph that is easier to implement on quantum hardware. The graph is generated such that each pair of nodes in the original hypergraph has a corresponding edge. Here, the edge $(0,1)$ has weight $w_{c}=(w_{a}+w_{b})/2$, whereas the other edges preserve the weights of the single hyperedge they correspondingly belong to. (d) Circuit implementation of the resulting graph with layers of SWAP gates. To limit the effect of hardware noise, we skip implementing the last two SWAP layers, at the cost losing some correlation with the original problem.
The second method consists of generating an Ansatz circuit with a potentially fully quadratic Hamiltonian whose parameters $\bm{\theta}$ are simultaneously optimized with $(\bm{\beta},\bm{\gamma})$ to minimize the energy of $H_{C}$. In particular, we perform the joint minimization
$\min_{\bm{\theta},\bm{\beta},\bm{\gamma}}\braket{\psi^{\prime}_{2}(\bm{\theta},\bm{\beta},\bm{\gamma})|H_{C}|\psi^{\prime}_{2}(\bm{\theta},\bm{\beta},\bm{\gamma})},$ (10)
where the Ansatz
[…]
is generated by the Hamiltonian
$H_{2}^{\prime}(\bm{\theta})=\sum_{i>j}\theta_{ij}Z_{i}Z_{j}+\sum_{i=0}^{n-1}\theta_{i}Z_{i}.$ (12)
The parameters $\bm{\theta}=\{\theta_{ij},\theta_{i}\}$ are optimized at the same time as the QAOA angles $\bm{\beta}$ and $\bm{\gamma}$. This differs from other Ansätze such as the multi-angle Ansatz (ma-QAOA) [herrman2021multiangle]. In ma-QAOA the Hamiltonian generating the Ansatz is built from the exact same terms as the cost operator $H_{C}$ but each Pauli term $k$ is associated with an independent QAOA angle $\gamma_{q,k}$ at each QAOA layer. Therefore, our $H_{2}^{\prime}(\bm{\theta})$ defines a quadratization of the original quartic Hamiltonian $H_{C}$.
@@
These quadratizations introduce an inconsistency between the cost operator and the operator that creates the Ansatz. Therefore, the convergence guarantee of a typical QAOA in the limit $p\to\infty$ is lost. Instead, the approach is inherently heuristic as it is based on an Ansatz circuit that retains only some of the structure of $H_{C}$, but which considerably reduces the circuit depth. The exact depth reduction depends on the density of the original problem. In the case of $H_{2}^{\prime}(\bm{\theta})$ the convergence as $p\to\infty$ may still persist as long as $H_{2}^{\prime}(\bm{\theta})$ shares the same ground state as $H_{C}$. The largest gate count reduction occurs for dense Hamiltonians, such as $H_{4}^{\text{full}}$, for which the reduction scales as $O(n^{2})$, see Fig. 1. For LABS, which has $O(n^{3})$ terms, the reduction factor is $O(n)$.
III.2 Expected performance of the quadratizations
We now benchmark the accuracy of the quadratizations by investigating depth-one and depth-two QAOA. We perform this assessment on the LABS Hamiltonians with 12 decision variables.
The depth-one QAOA angles $(\beta_{1}^{\star},\gamma_{1}^{\star})$ are trained with a coarse 2D grid scan with 15 points in the interval $0,\pi/2]$ followed by a Scipy optimization with COBYLA to refine the QAOA angles of the grid search. The depth-two angles are found with COBYLA from the transition state $(\beta_{1}^{\star},0,\gamma_{1}^{\star},0)$ as initial point [sack2023recursive]. The parameters $\bm{\theta}$ in $H_{2}^{\prime}(\bm{\theta})$ and the depth-$p$ QAOA angles are trained simultaneously with Scipy minimize and COBYLA to minimize the energy of $H_{C}$, see Fig. 3. The parameters in the clique expansion are optimized with Scipy and COBYLA independently from the QAOA parameters and converge to a low value of $f_{2}(\bm{\omega})$, see Fig. 3.
The optimized quantum circuits are sampled with the state-vector simulator of Qiskit. These samples are then converted to a cumulative distribution function of the energy of the LABS Hamiltonian $H_{C}$. We compare the clique expansion, the $H_{2}^{\prime}(\bm{\theta})$ projection, standard QAOA, and uniform sampling. The performance of the clique expansion is very close to random samples, see Fig. 4. Both standard depth-one and two QAOA outperform random sampling and the clique expansion, see dashed curves in Fig. 4. The depth-two standard QAOA, for example, samples the optimal -28 energy bit-string with a probability of 0.8%. The projection $H_{2}^{\prime}(\bm{\theta})$ does not find the optimal bit-string but produces good, yet sub-optimal, solution with high probability. For example, the depth-two QAOA with the $H_{2}^{\prime}(\bm{\theta})$ projection produces samples with a 97.1% approximation ratio, defined by $(E(x)-E_{\text{max}})/(E_{\text{min}}-E_{\text{max}})$, with a probability of 95.7%. Here, $E(x)$, $E_{\text{min}}$, and $E_{\text{max}}$ are the LABS energies of the candidate solution $x$, the lowest, and highest energies, respectively. The quadratization yields an interesting tradeoff since it only approximately solves the 12 node LABS problem but with a significantly shorter circuit depth. Indeed, the two-qubit gate depth and count of the Ansatz circuits based on the quadratization are 68 and 374 when transpiled with a SWAP network on a line of qubits [weidenfeller2202scaling]. By contrast, transpiling the standard QAOA Ansatz with the Qiskit transpiler at optimization level three on a fully connected network of qubits creates circuits with a two-qubit gate count and depth of 600 and 417, respectively. The increased gate count will likely worsen when the circuit is transpiled on a line of qubits. Irrespective of the qubit topology, this example shows the saving in quantum computational resources that the quadratization yields.
Figure 3: Convergence of the quadratizations as function of COBYLA iterations. The clique expansion, i.e., the minimization of Eq. (7), is shown as the dotted pink curve on the right $y$-axis. The solid blue and violet curves show the minimization of Eq. (10) for QAOA depths one and two, respectively.
Figure 4: Cumulative distribution function of the LABS Hamiltonian energy over $2^{14}$ samples obtained from different quantum circuits. The vertical solid line shows the minimum energy found by direct diagonalization. The dotted vertical line shows the best solution sampled from the $H_{2}^{\prime}(\bm{\theta})$ quadratization. The full energy range of the 12 qubits LABS problem is [28,110][-28,110].
The clique-based quadratization fails to produce good samples because it does not preserve the ground state of the problem. For example, the clique quadratization of the quartic term $Z_{3}Z_{2}Z_{1}Z_{0}$ produces the fully connected Hamiltonian $\sum_{i<j}^{3}Z_{i}Z_{j}$. The ground state of $Z_{3}Z_{2}Z_{1}Z_{0}$ is made of bit-strings with Hamming weight 1 or 3. By contrast, the ground state of the fully-connected quadratic Hamiltonian have Hamming weight 2. This also helps explain the success of the quadratization in Eq. (10) which constructs an Ansatz that minimizes the energy of the cost Hamiltonian.
@@ III.3 Noisy simulations
Figure 5: Evolution of expected performance of depth-two QAOA for LABS versus strength of depolarizing noise for a system of 12 qubits, using the standard Ansatz transpiled on hardware with full connectivity (pink) and with line connectivity (blue) compared to using the quadratized Ansatz transpiled on a line of qubits (purple). The circuit corresponding to the pink line has 600 two-qubit gates, the blue 2094, and the purple 374. Each data point is the average of 10 circuit runs and the error bars represent shot noise over 10 000 samples.
We now study the performance of the $H_{2}^{\prime}(\bm{\theta})$ quadratization and standard depth-two QAOA by setting up a noisy simulation in which we apply a depolarizing channel with strength $\lambda$ after each two-qubit gate. In Fig. 5, we show how the energy averaged over the best fraction $\alpha\in\{0.01,0.05.0.1\}$ of all samples evolves with $\lambda$. By varying $\alpha$ we see how the noise impacts the tail of the distribution of samples produced by the quantum circuit.
The quadratized Ansatz is less sensitive to noise, making it a viable option for noisy quantum computers. Generally, as $\lambda$ increases, the average energies of the best $\alpha$-fraction of samples increase, corresponding to a worse performance, as expected. Importantly, at $\alpha=0.01$ the Ansatz based on the QUBO projection (purple) produces samples that are statistically insensitive to the noise. By contrast, the energy of the QAOA Ansatze on a line of qubits (blue) and all-to-all connected qubits (pink) show a strong noise dependence, see Fig. 5(a). At $\alpha=0.1$, the QUBO Ansatz is $25\%$ and $36\%$ less sensitive [^2] Noise sensitivity is computed as $\frac{\Big\langle E_{\text{QUBO}}(\lambda=0.01)\Big\rangle-\Big\langle E_{\text{QUBO}}(\lambda=0.001)\Big\rangle}{\Big\langle E_{\text{LABS}}(\lambda=0.01)\Big\rangle-\Big\langle E_{\text{LABS}}(\lambda=0.001)\Big\rangle}$ than the other two types of Ansatze, respectively, see Fig. 5(c). These results indicate that the samples produced from a quadratized Ansatz are less sensitive to noise at the scale and noise strengths considered here.
The effectiveness of quadratizing the Ansatz depends on the noise strength and set $\alpha$ values. At $\alpha=0.01$ and noise strength $\lambda=0.001$, the standard QAOA Ansatz outperforms the quadratization. However, this advantage vanishes at $\lambda=0.004$ on a line connectivity. When considering the $5\%$ best samples, the quadratization is better than the standard QAOA for both all-to-all and line connectivity, regardless of noise strength, see Fig. 5(b).
These data show the benefits of quadratizing the Ansatz when working with hardware with small to intermediate noise relative to the size of the optimization problem. Furthermore, they reinforce the trade-off discussed in Fig. 4, namely that the quadratized Ansatz can increase the frequency of sampling low-energy, yet sub-optimal, states.
@@ IV Hardware-native circuits for QUBOs
Figure 6: Performance of QAOA as a function of depth and the number of SWAP layers. Each data point is the average of three RR3 graphs that require nine SWAP layers. The error bars show the standard deviation. (a) Approximation ratio of QAOA computed with a MPS evaluation of the energy. (b) Measured approximation ratios from 50 000 samples from ibm_fez. The horizontal dashed line is the $58.3\%$ average approximation ratio reported in Ref. [sack2024blackline]. (c) Value of the CVaR quantile $\alpha$ required to make the approximation ratio of the ibm_fez samples in (b) match those of the MPS energy evaluation in (a). The dashed lines show a simple theoretical expectation $\alpha_{\text{th}}$ based on the two-qubit gate fidelity.
We now study an approximate hardware-friendly compilation method of Ansatz circuits for QUBOs. The standard approach to route an arbitrary QAOA circuit utilizes a predetermined network of SWAP gates tailored to the qubit connectivity [weidenfeller2202scaling, matsuo2023sat, maciejewski2024design, Harrigan_2021]. We use a line of qubits to benefit from two-qubit gate cancellations, even though the IBM quantum hardware connectivity has a heavy-hex structure [weidenfeller2202scaling]. To fully implement $e^{-i\gamma H_{C}}$, we apply $k_{\text{max}}$ alternating layers of even and odd SWAP layers [^3] A SWAP layer is a depth-one circuit which applies SWAP gates to pairs of neighboring qubits, and it is called even if the first pair is $(0,1)$ and odd if it is $(1,2)$., see Fig. 2(d). Full connectivity is reached after n2n-2 SWAP layers. The SWAP network thus allows us to derive an approximate Ansatz by including only $k<k_{\text{max}}$ layers when implementing $e^{-i\gamma H_{C}}$. This allows us to implement only the time evolution of a subset $H_{C}(k)$ of the terms in $H_{C}$. The truncated $H_{C}(k)$ is based on the subgraph $G_{k}=(E_{k},V)\subseteq G=(E,V)$. Details on the transpilation are given in App. B. Importantly, the size $|E_{k}|$ depends on the initial mapping of the decision variables to the qubits. Following Ref. [matsuo2023sat] we optimize this initial mapping with a SAT solver to minimize the number of SWAP layers needed to implement the cost operator. Once the Ansatz is created, we optimize the QAOA parameters by minimizing the energy
[…]
@@
[…]
Our approach is similar to the time-block Ansatz design presented in Ref. [maciejewski2024design]. The main distinction is that a time-block Ansatz assigns different, independently optimized parameters to the edges of $H_{C}(k)$. Concretely, they replace the values $\gamma_{i,j}$ with corresponding vectors $\{\gamma_{i,j}^{e},\forall e\in E\}$.
We benchmark the truncated Ansatz on Max-Cut problems [karp1975computational] for all $k\in[0,k_{\max}]$. We thus aim to split the set $V$ of $n$ nodes of a graph $G=(E,V)$ in two parts such that the sum of the edges connecting the two subsets is maximized. Formally, a variable $x_{i}$ is assigned to each node $i\in V$ to indicate which side of the cut the node is on. Solving Max-Cut is then equivalent to finding the binary string $x$ that maximizes
$f(x)=\sum_{(i,j)\in E}\Big(x_{i}(1-x_{j})+x_{j}(1-x_{i})\Big).$ (15)
Mapping $f(x)$ to a Hamiltonian [lucas2014ising] results in the cost operator
$H_{\text{Max-Cut}}=\frac{1}{4}\sum_{(i,j)\in E}Z_{i}Z_{j},$ (16)